OpenStudy (iwanttogotostanford):

THREE QUESTIONS HELP PLEASE THREE Q"s:

OpenStudy (iwanttogotostanford):

3.) An electric field has a positive test charge of 5.00 E-5 C placed in it. The force on the test charge is 0.600 N. The magnitude of the electric field at the location of the test charge is 3.00 E5 N/C 3.00 E4 N/C 1.02 E4 N/C 120. E4 N/C 1.20 E4 N/C

OpenStudy (iwanttogotostanford):

@supersmart1001

OpenStudy (iwanttogotostanford):

@Qwertty123 @Vocaloid

OpenStudy (kky10997):

ans of 1st is binding energy

OpenStudy (kky10997):

ans of 2nd=option a

OpenStudy (supersmart1001):

r u guessing

OpenStudy (supersmart1001):

@kky10997 R U GUESSING

OpenStudy (kky10997):

no i am not

OpenStudy (supersmart1001):

how did u get those then

OpenStudy (kky10997):

in 1st proten's has negative charge so they repule is eachother so we need binding force to bind in a nucles and for that we need binding energy and that come from decreament from mass

OpenStudy (kky10997):

e=mc^2 we also know that so E_b = (n_Nm_N + n_Zm_Z - M)c^2,

OpenStudy (kky10997):

for 2nd let assume 1st we have n resistors and R1= eq resistance so 1/R1=1/r1+1/r2+....+1/rn now we increase no. resistors and we have N resistors N>n and now re resistance=R2 1/R2=1/r1+1/r2+.......+1/rn+1/rn+1+......+1/rN 1/R2=1/R1+1/rn+1+....+1/rN so 1/R2>1/R1 SO R2<R1

OpenStudy (kky10997):

@supersmart1001 now u understand

OpenStudy (supersmart1001):

idk what to do im totally lost too

OpenStudy (supersmart1001):

yeah

OpenStudy (iwanttogotostanford):

@xapproachesinfinity @

OpenStudy (iwanttogotostanford):

@welshfella @831CHEYENNE831

OpenStudy (kky10997):

srry chare on proton is +ve not -ve so sorry for that